Use the t-distribution to find a confidence interval for a difference in means μ 1 - μ 2 given the relevant sample results. Give the best estimate for μ 1 - μ 2 , the margin of error, and the confidence interval. Assume the results come from random samples from populations that are approximately normally distributed.

A 90% confidence interval for μ 1 - μ 2 using the sample results x ¯ 1 = 10.0 , s 1 = 2.2 , n 1 = 50 and x ¯ 2 = 13.1 , s 2 = 5.3 , n 2 = 50

Enter the exact answer for the best estimate and round your answers for the margin of error and the confidence interval to two decimal places.
Answer #1

Sample Mean 1 (X1) = 10 Sample Standard Deviation 1 (si) - 2.2 Sample Size 1 (N1) = 50 Sample Mean 2 (X2) = 13.1 Sample Stand

best estimate is 10-13.1=-3.1

margin of error = 1.668*0.812 = 1.3544

please like ??